# On Thermalization in A Nonlinear Variant of the Discrete NLS Equation

### v1

@@ A.1 Infinite temperature, $\beta\rightarrow 0$
$Z=(2\pi)^{N}\lambda_{0}^{N},\quad\frac{\partial Z}{\partial\lambda_{0}}=(2\pi)^{N}N\lambda_{0}^{N-1},$ (30)
$\frac{\partial\ln{Z}}{\partial\alpha}=\frac{1}{Z}\frac{\partial Z}{\partial\alpha}=\frac{1}{Z}\frac{\partial Z}{\partial\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\alpha}$
$=\frac{N}{(2\pi)^{N}\lambda_{0}^{N}}(2\pi)^{N}\lambda_{0}^{N-1}\frac{\partial\lambda_{0}}{\partial\alpha}=\frac{N}{\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\alpha}.$ (31)
$h$ $=-\frac{1}{N}\frac{\partial\ln{Z}}{\partial\beta}+\frac{1}{N}\frac{\alpha}{\beta}\frac{\partial}{\partial\alpha}\ln{Z}$
$=-\frac{1}{N}\frac{\partial\ln{Z}}{\partial\beta}-\alpha a=-\frac{1}{\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\beta}-\alpha a,$ (32)
where
$\frac{\partial\ln{Z}}{\partial\beta}=\frac{1}{Z}\frac{\partial Z}{\partial\beta}=\frac{1}{Z}\frac{\partial Z}{\partial\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\beta}$
$=\frac{1}{(2\pi)^{N}\lambda_{0}^{N}}(2\pi)^{N}N\lambda_{0}^{N-1}\frac{\partial\lambda_{0}}{\partial\beta}=\frac{N}{\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\beta}.$ (33)
Recalling the partition function of Eq. (II), we write
$Z$ $=(2\pi)^{N}\int_{0}^{\infty}\prod_{m=1}^{N}dA_{m}I_{0}\left[\frac{\beta D}{2}A_{m}A_{m+1}\right]$
$\times\exp{[-\beta\sum_{\ell}\frac{1}{8}(A_{\ell}^{2}+A_{\ell+1}^{2})+\frac{\alpha}{2}(A_{\ell}+A_{\ell+1})]}$
$=(2\pi)^{N}\int_{0}^{\infty}\prod_{m=1}^{N}dA_{m}I_{0}\left[\frac{\beta D}{2}A_{m}A_{m+1}\right]$
$\times\exp{[-\beta\sum_{\ell}\frac{1}{4}|A_{\ell}|^{2}+\alpha A_{\ell}]}$
$=(2\pi)^{N}\Bigg\{\int_{0}^{\infty}e^{-\beta\alpha A_{m}}\left[1-\frac{\beta A_{m}^{2}}{4}+\frac{1}{2}\left(\frac{\beta A_{m}^{2}}{4}\right)^{2}\right]$
$\times I_{0}\left[\frac{\beta D}{2}A_{m}A_{m+1}\right]dA_{m}\Bigg\}^{N}$
$=(2\pi)^{N}\!\left\{\!\int_{0}^{\infty}\!\!e^{-\beta\alpha A_{m}}\!\left[1\!-\!\frac{\beta A_{m}^{2}}{4}\!+\!\frac{1}{2}\!\left(\frac{\beta A_{m}^{2}}{4}\right)^{2}\right]\!dA_{m}\!\right\}^{N}$
$=(2\pi)^{N}I(\beta,\alpha)^{N},$ (34)
in which we used $I_{0}\left[\frac{\beta D}{2}A_{m}A_{m+1}\right]^{N}=1$ for $\beta=0$, and
$e^{-\beta(\frac{A^{2}}{4}+\alpha A)}=e^{-\beta\alpha A}e^{-\beta A^{2}/4}$
$=e^{-\beta\alpha A}\left(1-\frac{\beta A^{2}}{4}+\frac{1}{2}\left(\frac{\beta A^{2}}{4}\right)^{2}-\dots\right),$ (35)
$\prod_{m}e^{-\beta\alpha A_{m}}e^{-\beta A_{m}^{2}/4}=\prod_{m}e^{-\beta\alpha A_{m}}\left(1-\frac{\beta A_{m}^{2}}{4}+\dots\right).$ (36)
We use the gamma function $\Gamma(n)=(n-1)!$
$\frac{\Gamma(m+1)}{g^{m+1}}=\int_{0}^{\infty}x^{m}e^{-gx}dx,$ (37)
$\int x^{2}e^{-gx}dx=\Gamma(3)/(\beta\alpha)^{3},$ (38)
$\int x^{4}e^{-gx}dx=\Gamma(5)/g^{5}.$ (39)
To calculate the integral expression $I(\beta,\alpha)$ in Eq.(34):
$I=$ $\int_{0}^{\infty}e^{-\beta\alpha A_{m}}\left(1-\frac{\beta A_{m}^{2}}{4}+\frac{1}{2}\left(\frac{\beta A_{m}^{2}}{4}\right)^{2}-\dots\right)dA_{m}$
$=$ $\int_{0}^{\infty}e^{-\beta\alpha A_{m}}dA_{m}-\frac{\beta}{4}\int_{0}^{\infty}A_{m}^{2}e^{-\beta\alpha A_{m}}dA_{m}$
$+\frac{1}{2}\frac{\beta^{2}}{16}\int_{0}^{\infty}A_{m}^{4}e^{-\beta\alpha A_{m}}dA_{m}$
$=$ $-\frac{1}{\beta\alpha}e^{-\beta\alpha A_{m}}(0,\infty)-\frac{\beta}{4}\frac{\Gamma(3)}{(\beta\alpha)^{3}}+\frac{1}{2}\frac{\beta^{2}}{16}\frac{\Gamma(5)}{(\beta\alpha)^{5}}$
$=$ $\frac{1}{\beta\alpha}-\frac{\beta}{4}\frac{2}{(\beta\alpha)^{3}}+\frac{1}{2}\frac{\beta^{2}}{16}\frac{4.3.2}{(\beta\alpha)^{5}}$ (40)
For $\beta=0,I_{0}(0)=1$, then we can simply write $Z$ as
$Z$ $=(2\pi)^{N}I^{N}=(2\pi)^{N}\left[\frac{1}{\beta\alpha}-\frac{\beta}{2(\beta\alpha)^{3}}+\frac{3}{4}\frac{\beta^{2}}{(\beta\alpha)^{5}}\dots\right]^{N}$
$\approx(2\pi)^{N}\left(\frac{1}{\beta\alpha}-\frac{\beta}{2(\beta\alpha)^{3}}\right)^{N}$
$=(2\pi)^{N}\left(\frac{1}{\beta\alpha}\right)^{N}\left(1-\frac{1}{2}\frac{\beta}{(\beta\alpha)^{2}}\right)^{N}.$ (41)
in which $\beta\rightarrow 0,\alpha\rightarrow\infty,\beta\alpha={\rm const}$. We now take the logarithm to calculate the zero temperature $(a,h)$:
$ln(1-x)=-\sum_{n=1}^{\infty}\frac{x^{n}}{n}\approx-x-x^{2}/2-x^{3}/3\dots\approx-x-x^{2}/2$ (42)
Thus
$\ln(1-\frac{\beta}{2(\beta\alpha)^{2}})=-\frac{1}{2\beta\alpha^{2}}-\frac{1}{8\beta^{2}\alpha^{4}}.$ (43)
$\ln{Z}$ $=N\ln(2\pi)-N\ln(\beta\alpha)+Nln(1-\frac{\beta}{2(\beta\alpha)^{2}})$
$\approx N\ln(2\pi)-N\ln(\beta\alpha)-\frac{N}{2\beta\alpha^{2}}-\frac{N}{8\beta^{2}\alpha^{4}}$ (44)
$\frac{\partial\ln{Z}}{\partial\alpha}=-\frac{N}{\beta\alpha}\beta-\frac{N}{2\beta}(-2)\alpha^{-3}=-\frac{N}{\alpha}+\frac{N}{\beta\alpha^{3}}-\frac{N}{2\beta^{2}\alpha^{5}}$ (45)
$\frac{\partial\ln{Z}}{\partial\beta}=-\frac{N}{\beta}+\frac{N}{2(\beta\alpha)^{2}}+\frac{\beta N}{4(\beta\alpha)^{4}}$ (46)
Recalling Eq. (10)
$a$ $=-\frac{1}{\beta N}\frac{\partial\ln{Z}}{\partial\alpha}=-\frac{1}{\beta N}\left(-\frac{N}{\alpha}+\frac{N}{\beta\alpha^{3}}-\frac{N}{2\beta^{2}\alpha^{5}}\right)$
$=\frac{1}{\alpha\beta}-\frac{1}{\beta^{2}\alpha^{3}}+\frac{1}{2\beta^{3}\alpha^{5}}\underset{\beta\rightarrow 0}{\approx}\frac{1}{\alpha\beta}$ (47)
Recalling Eq. (A.1)
$h$ $=-\frac{1}{N}\frac{\partial\ln{Z}}{\partial\beta}-\alpha a$
$=\left(\frac{1}{\beta}-\frac{1}{2\alpha^{2}\beta^{2}}-\frac{1}{4\beta^{3}\alpha^{4}}\right)-\alpha\left(\frac{1}{\alpha\beta}-\frac{1}{\beta^{2}\alpha^{3}}+\frac{2}{4\beta^{3}\alpha^{5}}\right)$
$=\frac{1}{2\alpha^{2}\beta^{2}}-\frac{3\beta}{4(\beta\alpha)^{4}}\underset{\beta=0}{\approx}\frac{1}{2\alpha^{2}\beta^{2}}=\frac{a^{2}}{2}$ (48)
in which we used
$\ln{Z}$ $\approx Nln(2\pi)-N\ln(\beta\alpha)-\frac{N}{2\alpha^{2}\beta}$ (49)
$\frac{\partial\ln{Z}}{\partial\beta}$ $=-\frac{N}{\beta}+\frac{N}{2\alpha^{2}\beta^{2}}.$ (50)
We found $h=a^{2}/2$ in Eq. (48) for $\beta\rightarrow 0$.
A.2 Zero temperature ($\beta\rightarrow\infty$)
When the temperature is zero, $\beta\rightarrow\infty$, the minimum energy is realized when all phases keep the same phase difference $\pi$ such that $\theta_{\ell}=\ell\pi$, and $A_{\ell}=a$. This leads to
$b_{\ell}=\sqrt{A_{\ell}}e^{i\theta_{\ell}}=\sqrt{a}e^{i\ell\pi},\quad b_{\ell}^{2}=ae^{2i\ell\pi},$ (51)
${b_{\ell}^{*}}^{2}=ae^{-2i\ell\pi}\quad|b_{\ell}|^{4}=a^{2}$ (52)
${b_{\ell}^{*}}^{2}{b_{\ell-1}}^{2}+{b_{\ell}}^{2}{b_{\ell-1}^{*}}^{2}=a^{2}e^{-2i\ell\pi+2i(\ell-1)\pi}+a^{2}e^{2i\ell\pi-2i(\ell-1)\pi}$
$=a^{2}(e^{-2i\pi}+e^{2i\pi})=2a^{2}\cos{(2\pi)}=2a^{2}$ (53)
Inserting Eq. (51) and Eq. (A.2) into the Hamiltonian Eq. (3) gives
$h$ $=\frac{H}{N}=\frac{1}{N}\sum_{\ell}\frac{1}{4}|b_{\ell}|^{4}-\frac{D}{4}({b_{\ell}^{*}}^{2}b_{\ell-1}^{2}+b_{\ell}^{2}{b_{\ell-1}^{*}}^{2})$
$=\frac{1}{N}N\left(\frac{a^{2}}{4}-\frac{D}{4}2a^{2}\right)=\frac{a^{2}}{4}(1-2D)=\frac{a^{2}}{4}(1-2D)$ (54)
Appendix B Staggered compacton solutions for m=1m=166
@@
$x=\frac{D^{2}+D-1}{D^{2}+8D-5}A,y=\frac{(2D-1)A}{D^{2}+8D-5},z=\frac{-D^{2}+2D-1}{D^{2}+8D-5}A,$
@@
$x=\frac{(D^{2}-1)A}{2(2D^{2}+2D-3)},y$ $=\frac{(D^{2}+D-1)A}{2(2D^{2}+2D-3)},$ (62)
$z=\frac{(D-1)A}{2(2D^{2}+2D-3)},$
@@
“—” denotes a singular coefficient due to a zero denominator.

### v2

@@ Appendix A Phase Diagram for Toy Model
$J(\beta,\mu):=\int_{0}^{\infty}e^{-\mu A}\,e^{-(\beta/4)A^{2}}\,dA.$ (30)
Expanding $e^{-(\beta/4)A^{2}}=1-(\beta/4)A^{2}+\mathcal{O}(\beta^{2})$ and using
$\int_{0}^{\infty}e^{-\mu A}\,dA=\frac{1}{\mu},\qquad\int_{0}^{\infty}A^{2}e^{-\mu A}\,dA=\frac{2}{\mu^{3}},$ (31)
we find
$J(\beta,\mu)=\frac{1}{\mu}-\frac{\beta}{2\mu^{3}}+\mathcal{O}(\beta^{2})=\frac{1}{\mu}\left(1-\frac{\beta}{2\mu^{2}}+\mathcal{O}(\beta^{2})\right).$ (32)
Hence,
$\ln Z$ $=N\ln(2\pi)+NJ(\beta,\mu)$ (33)
$=N\ln(2\pi)-N\ln(\beta\alpha)-\frac{N}{2\beta\alpha^{2}}+\mathcal{O}(\beta).$
Using the thermodynamic identities [Eq. (10)],
$a=-\frac{1}{\beta N}\frac{\partial\ln Z}{\partial\alpha},\qquad h=-\frac{1}{N}\frac{\partial\ln Z}{\partial\beta}-\alpha a,$ (34)
and differentiating Eq. (33) yields
$\frac{\partial\ln Z}{\partial\alpha}\!=\!-\frac{N}{\alpha}+\frac{N}{\beta\alpha^{3}}+\mathcal{O}(\beta),\quad\frac{\partial\ln Z}{\partial\beta}\!=\!-\frac{N}{\beta}+\frac{N}{2\beta^{2}\alpha^{2}}+\mathcal{O}(1).$ (35)
Therefore,
$a=\frac{1}{\beta\alpha}+\mathcal{O}(\beta),\qquad h=\frac{1}{2\beta^{2}\alpha^{2}}+\mathcal{O}(1),$ (36)
and eliminating $\beta\alpha$ gives the infinite-temperature curve
$h(\beta=0)=\frac{a^{2}}{2}.$ (37)
Zero temperature ($\beta\to\infty$)
In the zero-temperature limit ($\beta\to\infty$), the lowest-energy homogeneous-norm configuration ($A_{\ell}=a$) is obtained for a staggered phase pattern with $\Delta\theta=\pi$, i.e., $\theta_{\ell}=\ell\pi$. Then
$b_{\ell}=\sqrt{a}\,e^{i\ell\pi},\qquad|b_{\ell}|^{4}=a^{2},$ (38)
and
${b_{\ell}^{*}}^{2}b_{\ell-1}^{2}+b_{\ell}^{2}{b_{\ell-1}^{*}}^{2}$ $=2a^{2}\cos\!\big(2(\theta_{\ell}-\theta_{\ell-1})\big)$ (39)
$=2a^{2}\cos(2\pi)=2a^{2}.$
Substituting Eqs. (38)–(39) into Eq. (3) yields the corresponding energy density
$h=\frac{H}{N}=\frac{a^{2}}{4}-\frac{D}{4}\,2a^{2}=\frac{a^{2}}{4}(1-2D).$ (40)
Appendix B Staggered $m$-site compactons: $m=1,\dots,6$
@@
$x$ $=\frac{D^{2}+D-1}{D^{2}+8D-5}\,A,$ $y$ $=\frac{2D-1}{D^{2}+8D-5}\,A,$
$z$ $=\frac{-D^{2}+2D-1}{D^{2}+8D-5}\,A.$ (46)
@@
$x=\frac{(D^{2}-1)A}{2(2D^{2}+2D-3)},\qquad y$ $=\frac{(D^{2}+D-1)A}{2(2D^{2}+2D-3)},$
$z=\frac{(D-1)A}{2(2D^{2}+2D-3)},$ (49)
@@
“—” denotes a singular coefficient (zero denominator).
Negative entries indicate non-admissible amplitudes.